OpenStudy (jojo_rawrs):

Solve this quadratic equation using the quadratic formula. 2x2 - 10x + 7 = 0

pooja195 (pooja195):

List your abc values first

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

pooja195 (pooja195):

\[\huge~\rm~x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

JoJo you here?

OpenStudy (jojo_rawrs):

yup

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

So, you need to take out what the A, B And C values are

OpenStudy (jojo_rawrs):

omg i hate math no joke

OpenStudy (anonymous):

Lol I'm in algebra 1 too.

OpenStudy (anonymous):

It CAN be fun though

OpenStudy (jojo_rawrs):

this sucks

OpenStudy (anonymous):

Alright, so your A = 2, B = -10 and C = 7

OpenStudy (anonymous):

Now plug those into the equation ppl gave u up top.

OpenStudy (jojo_rawrs):

i try it

OpenStudy (anonymous):

Show me the plugged in equation.

OpenStudy (anonymous):

Hello?

OpenStudy (jojo_rawrs):

yes sorry i was doing something

OpenStudy (anonymous):

\[x = \frac{ -(-10)\pm \sqrt{-10^2 - 4(2)(7)} }{ 2(2)}\]

OpenStudy (anonymous):

Can you solve that?

OpenStudy (jojo_rawrs):

no to be honest

OpenStudy (anonymous):

Did you even go through the lesson? :3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Do you need help solving it?